A 3kg object travels in a uniform circle at a constant speed of 12m/s. The centripetal force acting on the object is 110 Newtons.


Calculate the time required for the the object to complete one full circle.


1
Expert's answer
2022-01-30T13:52:17-0500

Given:

"m=3\: \rm kg"

"v=12\: \rm m/s"

"F_c=110\:\rm N"


The cetripetal force

"F_c=\frac{mv^2}{r}"

So, the radius of the circle

"r=\frac{mv^2}{F_c}=\frac{3*12^2}{110}=3.93\: \rm m"

The period of rotation

"T=\frac{2\pi r}{v}=\frac{6.28*3.93}{12}=2.1\:\rm s"